18. A box of animal crackers contains five hippos, two lions, three zebras, and four elephants. Find the probability if one animal cracker is chosen at random. a. Pla hippo b. P(not an elephanf) c. P(an elephant or a lion) lu 19. Suppose you have a dark closet containing seven blue shirts, five yellow shirts, and eight white shirts. You pick two shirts at random from the closet. Find each probability. a. P(blue then yellow) with replacing b. P(blue then yellow) without replacing Id Plyellow then yellow) with replacing d. P(yellow then yellow) without replacing e. e. P(yellow then white) with replacing f. P(yellow then white) without replacing g. P(blue then blue) with replacing h. P(blue then blue) without replacing 20. Suppose you draw two tennis balls at random from a bag containing seven pink, four white, three yellow, and two striped balls. Find each probability. a. P(yellow then pink) with replacing b. P(yellow then pink) without replacing P(pink then pink) with replacing d. P(pink then pink) without replacing c. e. P(striped then striped) with replacing . P(striped then striped) without replacing E Pipink then white with replacing h. P(pink then white) without replacing
Answer #1

Please dont hesitate to give a 'thumbs up' for the answer in case you're satisfied with the answer.

18.

5 hippos, 2 lions, 3 zebras, and 4 elephants

a.P(a hippo) = no. of total hippos/total animals = 5/(5+2+3+4) = 5/14

b.P(not elephant) = no. of animals not elephants/total animals = (14-4)/14 = 10/14

c.P(an elephant or a lion) = 4+2/14 = 6/14 = 3/7